The daily cost to a printer is modeled by the function C(x)=3x+1800 where C(x) is the total daily cost and x is the number of copies printed. Find C(4000), and interpret its meaning.
C(4000)=
This value represents
The number of copies manufactured when the daily cost is 4000
The product of the daily cost and the number of copies
The daily cost when 4000 copies are printed

Answer by ikleyn(52781)

.
The daily cost to a printer is modeled by the function C(x)=3x+1800 where C(x) is the total daily cost and x is the number of copies printed. Find C(4000), and interpret its meaning.
C(4000)=
This value represents
(a) The number of copies manufactured when the daily cost is 4000
(b) The product of the daily cost and the number of copies
(c) The daily cost when 4000 copies are printed
~~~~~~~~~~~~~~~~~~~~~~~~

C(4000) = 3*4000+1800 = 12000+1800 = 13800.

You simply should substitute the value of 4000 instead of x into the formula
and make these simple calculations.

For the other part of the problem, the answer is option (c).

It is so simple that everyone who is able to read and understand the text,
can not make an error.

Solved.

This problem is of the complexity level 2 x 2 = 4.